OpenStudy (anonymous):

Given f(x) = -x + 15, what is f(-3) ? A. –45 B. 12 C. 18 D. 45

OpenStudy (solomonzelman):

plug in -3 for x. Like if I was to solve for f(-5) this is what I would do. \[f(x) = -x + 15\]\[f(-5) = -(-5) + 15\]\[f(-5) =+5 + 15\]\[f(-5) =20\]

OpenStudy (anonymous):

okay. I'll try it

OpenStudy (solomonzelman):

Yes, and tell me what you get for f(-3) after you solve it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 18. Is that right?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes , way to do it !!

OpenStudy (solomonzelman):

Can I test you with similar question ?

OpenStudy (anonymous):

Um okay

OpenStudy (solomonzelman):

Given f(x) =x^2+4 . Find f(-1) ?

OpenStudy (anonymous):

I'm not sure but maybe -1?